# Gapless symmetry protected topological phases and generalized deconfined critical points from gauging a finite subgroup

### v1

@@ III.2 Emergent mixed anomaly
In the low energy theory, there is an emergent mixed anomaly between the quotient $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$ (considering PBC for simplicty). One manifestation of the mixed anomaly is that $\tilde{U}(1)$ and $W$ cannot be simultaneously realized on-site. Here $\tilde{U}(1)$ and $W$ seem to be realized on-site. However, $\tilde{U}(1)$ is only exact when the gauge condition in Eq. (18) is enforced, i.e., $K\to\infty$, and the Hilbert space then is not a tensor product. We can follow Ref. [33] to eliminate the gauge constraints and find that either $\tilde{U}(1)$ or $W$ is realized non-on-site (see Appendix B).
If we turn on the background gauge field $A^{\tilde{U}(1)}$ and $A^{W}$ associated with $\tilde{U}(1)$ and $W$ respectively, the mixed anomaly is characterized by a response action in $(2+1)$-D: $\omega=A^{W}\cup dA^{\tilde{U}(1)}/4\pi$ where $\cup$ is a cup product, a discrete analogue of the wedge product of differential forms  [^3] The reader may treat it simply as a wedge product.. Thus, $\tilde{U}(1)$ and $W$ cannot be gauged consistently in the $(1+1)$-D system we are studying which may be viewed as the boundary of the $(2+1)$-D bulk (Appendix D). The system, viewed as the boundary of the bulk, usually cannot have boundaries because the boundary of a boundary is an empty set. This is not a contradiction because in our work the mixed anomaly is emergent in the low energy sector. After some suitable modifications of boundary terms, we can put this model on a lattice with open boundary conditions while preserving both symmetries in the Hamiltonian, which plays a crucial role in the “Higgs = SPT” argument in Refs. [27, 28] and our argument in the following for the gapless SPT phase.
As a result of the emergent mixed anomaly, the ground state cannot be trivially gapped, meaning that in a gapped phase $W$ must be spontaneously broken since $\tilde{U}(1)$ cannot be spontaneously broken due the Mermin-Wagner theorem. Indeed, $W$ is spontaneously broken in the insulating phase, while in the superfluid phase both $\tilde{U}(1)$ and $W$ are preserved. As we have discussed above, in the superfluid phase, both $\langle b^{\dagger}_{i}b^{\dagger}_{i}b_{j}b_{j}\rangle$ and $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|$ are shown to have a power law decay (Fig. 2(a) and 3(a)). These correlated ordering/disordering behaviors are already encoded in the relation $\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle=\langle\prod_{i\in R}(-1)^{n_{i}}\rangle=\langle\tilde{X}_{R}(2\pi)\rangle$ we mentioned earlier. This relation implies that $\tilde{U}(1)$ is preserved if $W$ is spontaneously broken and that $\tilde{U}(1)$ is spontaneously broken if $W$ is preserved.
@@ III.3 Gapless SPT phase
In this subsection, we show that the superfluid phase is a gapless SPT phase.
@@
Next, we discuss the finite size scaling in the bulk gap $\Delta_{\text{bulk}}$ to show the bulk is indeed gapless in the thermodynamic limit. As we can see from Fig. 1(d), the bulk gap, the first excited state from the doubly degenerate ground state, is inversely proportional to the system size $L$. In the thermodynamic limit, the bulk correlation length $\xi_{b}$ diverges and the bulk becomes gapless  [^4] Generically, if $\Delta_{\text{bdry}}$ also has a dependence on the system size, the edge degeneracy refers to an exponential decay or at least a decay faster than the bulk gap $\Delta_{\text{bulk}}$ [13, 14, 15, 18].. This is compatible with the fact that there is a mixed anomaly between $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. As we have already mentioned, both $\tilde{U}(1)$ and $W$ are preserved, which is supported by algebraically decaying $\langle b_{i}^{\dagger}b_{i}^{\dagger}b_{j}b_{j}\rangle$ (Fig. 2(a)) and $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|$ (Fig. 3(a)). The gap $\Delta_{\text{bulk}}$ in the insulator phase, on the other hand, remains finite in the thermodynamical limit, as extrapolated by finite-size scaling (Fig. 1(e)). Thus, we have showed that the superfluid phase is a gapless SPT phase.
@@
where $A^{W}$ and $A^{P}$ is the background field of $W$ and $P$ in the spacetime $M$, respectively. If $A^{P}$ and $A^{W}$ are flat, $dA^{P}=0$ and $A^{W}=0$. Then if $M$ is closed, $\alpha$ is gauge invariant and describes a SPT phase protected by $W$ and $P$. Indeed, if $M$ has a nontrivial boundary, there can be an open Wilson line terminating on the $\partial M$ where $dA^{W}\neq 0$. Then the action $S_{M}=2\pi\int_{M}\alpha$ changes by is $\lambda A^{W}/2$ under the gauge transformation $A^{P}\to A^{P}+d\lambda^{P}$. To compensate this change, there must be edge modes.
@@
Indeed, we have already seen that there is an emergent mixed anomaly between $\tilde{U}(1)$ and $W$. If we denote the $(2+1)$-D bulk as $Y$ such that its boundary is the $(1+1)$-D spacetime $M$ that we are studying, i.e. $\partial Y=M$, then the corresponding anomaly action can be written as $\omega=A^{W}\cup dA^{\tilde{U}(1)}/4\pi$ where $A^{\tilde{U}(1)}$ and $A^{W}$ are extended into $Y$. Note that $\alpha$ and $\omega$ satisfy the anomaly vanishing equation [18]
@@
Conceptually, constructing intrinsically gapless SPT phases using partial gauging as we discussed in this work is easier than the approach used in Ref. [18]. The mixed anomaly between the quotient symmetry and the dual symmetry is a direct consequence of the nontrivial group extension and does not depend on the details of the Hamiltonian. In Ref. [18], the authors considered a fermionic system, but the analysis above is obviously generalizable to fermionic systems even though the fermionic parity cannot be spontaneously broken and spin structures may need to be taken into account. As long as there is no other SSB order, the gauged Luttinger liquid of spinless fermions with $W=\prod_{i}\sigma_{i+1/2}^{z}$ preserved is a gapless SPT phase.
@@ III.4 Conformal field theory
This expectation again can be verified by the DMRG results. We first check the central charge $c$ in the superfluid phase in the gauged system. Indeed, as shown in Fig. 2(d), the subsystem entanglement entropy $S_{E}=-\text{tr}(\rho_{R}\ln\rho_{R})$, associated with the reduced density matrix $\rho_{R}$ of a subsystem $R$, scales linearly with the factor $\lambda(l)\equiv\frac{1}{6}\log\left(\frac{2L}{\pi}\sin\left(\frac{\pi l}{L}\right)\right)$ [48]. Here $L$ is the total system size of the open chain and $l$ is the size of the subsystem $R$ on one side of the entanglement cut. The slope gives us the central charge $c\sim 1$, the nominal central charge of a compact boson CFT.
@@
where $z=\exp(\tau+ix)$, and the free boson field $\phi$ is compactified on a circle of radius $R$, i.e. $\phi(z,\bar{z})\sim\phi(z,\bar{z})+2\pi R$. Split $\phi(z,\bar{z})$ into the left-moving and the right-moving components: $\phi(z,\bar{z})=X_{L}(z)+X_{R}(\bar{z})$. Then the local primary operators are:
@@ III.5 Perturbations
In the discussions above, we have argued that the superfluid is a gapless SPT is protected by $U(1)$ and $W$. As long as the Hamiltonian (and the boundary conditions) preserve $U(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$, the edge degeneracy is protected. For instance, as verified by DMRG, adding a term $\sum\sigma^{z}_{i-1/2}\sigma^{z}_{i+1/2}$ does not lift the degeneracy. Adding a small perturbation $\lambda\sum\sigma^{x}_{i+1/2}$ however breaks this symmetry. The edges open up a small gap and are no longer degenerate. The situation is different in higher dimensions when the protecting symmetries include higher-form symmetries. Breaking higher-form symmetries explicitly may not lift the edge degeneracy.
@@
Having considered the 1-D case, we can generalize the analysis to higher dimensions. In this section, we consider the emergent $\mathbb{Z}_{2}$-gauged Bose-Hubbard model on a 2-D square lattice (Fig 4(a)):
@@ IV.1 Emergent mixed anomaly
Similar to the 1-D case, the ground state of the system cannot be trivially gapped (i.e. nondegenerate, gapped, and symmetric under both symmetries). This consequence strongly constrains the phase diagram. As we will argue in this work, the critical point inherited from the ordinary superfluid-insulator transition now becomes a generalized DQCP between a gapless SPT phase where $\tilde{U}(1)$ is spontaneously broken and a SET phase where the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry is spontaneously broken (Fig. 4(b)).
@@ IV.2 Gapless SPT phase
To show that the state is SPT, we place it on lattice with open boundaries that preserve the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. We consider a half-plane geometry with an infinitely long boundary (Fig. 5(a)), where the “rough” boundary has dangling bonds sticking out so that properly modified local gauge constraints Eq. (18) are still well-defined [52]. Then similar to the 1-D case, we can argue that if $W$ is preserved, there will necessarily be a boundary degeneracy as follows. We choose the symmetry generator of $W=\prod_{e\in\gamma}\sigma_{e}^{z}$ to be a Wilson line with one end terminating on the boundary and the other end either extending to infinity or terminating on a boson creation/annihilation operator. With $P$ and $W$ preserved in the bulk, the anti-commutativity of $P$ and $W$ implies a ground state degeneracy, which in this case necessarily comes from the boundary. This defines a SPT phase. In fact, the Higgs condensate $\langle b_{i}^{\dagger}\sigma_{i+1/2}^{z}...\sigma_{j-1/2}^{z}b_{j}\rangle$ can be viewed as a SPT string order parameter here.
@@ IV.3 Excitations in the superfluid phase
We have argued that the superfluid phase is a gapless SPT phase. Here we discuss the excitations in the phase: vortices, Goldstone bosons, and domain walls. Since the $\mathbb{Z}_{2}$-gauging process amounts to projecting out the $\mathbb{Z}_{2}$-charged sector and adding the twisted sector to the theory and the process relates some order parameters to disorder parameters [33, 51], many physical properties of this new gapless SPT phase can be inferred directly from its superfluid parent state. As we have already seen above, the disorder parameter $|\langle\hat{X}_{R}(\theta)\rangle|$ for $\tilde{U}(1)$ is a (fractional) order parameter for the dual ^$\hat{\mathbb{Z}}_{2}$. Similarly, $|\langle b_{i}^{\dagger}(\prod_{e\in\gamma_{ij}}\sigma_{e}^{z})b_{j}\rangle|$ where $\gamma_{ij}$ connects sites $i$ and $j$ (Fig. 5(c)), viewed as a disorder parameter for the dual ^$\hat{\mathbb{Z}}_{2}$, serves as a (fractional) order parameter for $\tilde{U}(1)$.
To guarantee the exactness of the 1-form symmetry, we imposed the zero-flux condition $\prod_{e\in\partial p}\sigma_{e}^{z}=1$, which suppresses all (dynamical) $\pi$-vortices excitations. Equivalently, an open ’t Hooft line which would end on a pair of $\pi$-vortices (Fig. 5(c)) are also suppressed. In the Higgs phase, the condensate phase $\varphi$ is locked to the vortcies. The dual 1-form symmetry measures the $\mathbb{Z}_{2}$ winding of condensate phase $\varphi$, which takes values in $n\pi$ for integer $n$ modulo 2, along noncontractible cycles, which is equivalent to insertions of $\pi$-fluxes across noncontractible cycles (see Fig. 5(d)). We should compare the scenario with that in the “parent” superfluid phase before gauging where there is an emergent 1-form $U(1)$ symmetry in the low energy sector [22, 56, 57]. Charged objects of the emergent $U(1)$ are the winding of $\varphi$, taking values in $2n\pi$ for integer $n$. $2\pi$-vortices explicitly breaks this emergent dual $U(1)$ 1-form symmetry. However, since they are neutral under the dual ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry $W$ after the effective gauging, they do not destroy the exactness of the dual ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry.

### v2

@@ III.2 Emergent mixed anomaly
In the low energy theory, there is an emergent mixed anomaly between the quotient $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$ (considering PBC for simplicity). One manifestation of the mixed anomaly is that $\tilde{U}(1)$ and $W$ cannot be simultaneously realized on-site. Here $\tilde{U}(1)$ and $W$ seem to be realized on-site. However, $\tilde{U}(1)$ is only exact when the gauge condition in Eq. (18) is enforced, i.e., $K\to\infty$, and the Hilbert space then is not a tensor product. We can follow Ref. [33] to eliminate the gauge constraints and find that either $\tilde{U}(1)$ or $W$ is realized non-on-site (see Appendix B).
If we turn on the background gauge field $A^{\tilde{U}(1)}$ and $A^{W}$ associated with $\tilde{U}(1)$ and $W$ respectively, the mixed anomaly is characterized by a response action in $(2+1)$-D: $\omega=A^{W}\cup dA^{\tilde{U}(1)}/4\pi$ where $\cup$ is a cup product, a discrete analog of the wedge product of differential forms  [^3] The reader may treat it simply as a wedge product.. Thus, $\tilde{U}(1)$ and $W$ cannot be gauged consistently in the $(1+1)$-D system we are studying which may be viewed as the boundary of the $(2+1)$-D bulk (Appendix D). The system, viewed as the boundary of the bulk, usually cannot have boundaries because the boundary of a boundary is an empty set. This is not a contradiction because in our work the mixed anomaly is emergent in the low energy sector. After some suitable modifications of boundary terms, we can put this model on a lattice with open boundary conditions while preserving both symmetries in the Hamiltonian, which plays a crucial role in the “Higgs = SPT” argument in Refs. [27, 28] and our argument in the following for the gapless SPT phase.
As a result of the emergent mixed anomaly, the ground state cannot be trivially gapped, meaning that in a gapped phase $W$ must be spontaneously broken since $\tilde{U}(1)$ cannot be spontaneously broken due the Mermin-Wagner theorem. Indeed, $W$ is spontaneously broken in the insulating phase, while in the superfluid phase both $\tilde{U}(1)$ and $W$ are preserved. As we have discussed above, in the superfluid phase, both $\langle b^{\dagger}_{i}b^{\dagger}_{i}b_{j}b_{j}\rangle$ and $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|$ are shown to have a power law decay [Fig. 2(a) and 3(a)]. These correlated ordering/disordering behaviors are already encoded in the relation $\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle=\langle\prod_{i\in R}(-1)^{n_{i}}\rangle=\langle\tilde{X}_{R}(2\pi)\rangle$ we mentioned earlier. This relation implies that $\tilde{U}(1)$ is preserved if $W$ is spontaneously broken and that $\tilde{U}(1)$ is spontaneously broken if $W$ is preserved.
@@ III.3 Gapless SPT phase
In this section, we show that the superfluid phase is a gapless SPT phase.
@@
Next, we discuss the finite size scaling in the bulk gap $\Delta_{\text{bulk}}$ to show the bulk is indeed gapless in the thermodynamic limit. As we can see from Fig. 1(d), the bulk gap, the first excited state from the doubly degenerate ground state, is inversely proportional to the system size $L$. In the thermodynamic limit, the bulk correlation length $\xi_{b}$ diverges and the bulk becomes gapless  [^4] Generically, if $\Delta_{\text{bdry}}$ also has a dependence on the system size, the edge degeneracy refers to an exponential decay or at least a decay faster than the bulk gap $\Delta_{\text{bulk}}$ [13, 14, 15, 18].. This is compatible with the fact that there is a mixed anomaly between $\tilde{U}(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. As we have already mentioned, both $\tilde{U}(1)$ and $W$ are preserved, which is supported by algebraically decaying $\langle b_{i}^{\dagger}b_{i}^{\dagger}b_{j}b_{j}\rangle$ [Fig. 2(a)] and $|\langle\sigma_{i}^{x}\sigma_{j}^{x}\rangle|$ [Fig. 3(a)]. The gap $\Delta_{\text{bulk}}$ in the insulator phase, on the other hand, remains finite in the thermodynamical limit, as extrapolated by finite-size scaling [Fig. 1(e)]. Thus, we have showed that the superfluid phase is a gapless SPT phase.
@@
where $A^{W}$ and $A^{P}$ is the background field of $W$ and $P$ in the spacetime $M$, respectively. If $A^{P}$ and $A^{W}$ are flat, $dA^{P}=0$ and $dA^{W}=0$. Then if $M$ is closed, $\alpha$ is gauge invariant and describes an SPT phase protected by $W$ and $P$. Indeed, if $M$ has a nontrivial boundary, there can be an open Wilson line terminating on the $\partial M$ where $dA^{W}\neq 0$. Then the action $S_{M}=2\pi\int_{M}\alpha$ changes by is $\lambda A^{W}/2$ under the gauge transformation $A^{P}\to A^{P}+d\lambda^{P}$. To compensate this change, there must be edge modes.
@@
Indeed, we have already seen that there is an emergent mixed anomaly between $\tilde{U}(1)$ and $W$. If we denote the $(2+1)$-D bulk as $Y$ such that its boundary is the $(1+1)$-D spacetime $M$ that we are studying, i.e., $\partial Y=M$, then the corresponding anomaly action can be written as $\omega=A^{W}\cup dA^{\tilde{U}(1)}/4\pi$ where $A^{\tilde{U}(1)}$ and $A^{W}$ are extended into $Y$. Note that $\alpha$ and $\omega$ satisfy the anomaly vanishing equation [18]
@@
Conceptually, constructing intrinsically gapless SPT phases using partial gauging as we discussed in this work is easier than the approach used in Ref. [18]. The mixed anomaly between the quotient symmetry and the dual symmetry is a direct consequence of the nontrivial group extension and does not depend on the details of the Hamiltonian. In Ref. [18], the authors considered a fermionic system, but the analysis above is obviously generalizable to fermionic systems even though the fermionic parity cannot be spontaneously broken and spin structures may need to be taken into account. Indeed, a similar analysis can be found in Ref. [49]. There, they gauged the fermionic parity of a free spinless fermionic chain. Hubbard interactions can also be added. As long as there is no other SSB order, the gauged Luttinger liquid of spinless fermions with $W$ preserved is a gapless SPT phase. Spinful fermions can also be studied. Note that gauging the fermionic parity produces a bosonic theory that does not depend on spin structures. In particular, gauging the parity of a 1-D Dirac fermion theory yields the $XY$ model with a BKT transition [50].
@@ III.4 Conformal field theory
This expectation again can be verified by the DMRG results. We first check the central charge $c$ in the superfluid phase in the gauged system. Indeed, as shown in Fig. 2(d), the subsystem entanglement entropy $S_{E}=-\text{tr}(\rho_{R}\ln\rho_{R})$, associated with the reduced density matrix $\rho_{R}$ of a subsystem $R$, scales linearly with the factor $\lambda(l)\equiv\frac{1}{6}\ln\left(\frac{2L}{\pi}\sin\left(\frac{\pi l}{L}\right)\right)$ [52]. Here $L$ is the total system size of the open chain and $l$ is the size of the subsystem $R$ on one side of the entanglement cut. The slope gives us the central charge $c\sim 1$, the nominal central charge of a compact boson CFT.
@@
where $z=\exp(\tau+ix)$, and the free boson field $\phi$ is compactified on a circle of radius $R$, i.e., $\phi(z,\bar{z})\sim\phi(z,\bar{z})+2\pi R$. Split $\phi(z,\bar{z})$ into the left-moving and the right-moving components: $\phi(z,\bar{z})=X_{L}(z)+X_{R}(\bar{z})$. Then the local primary operators are:
@@ III.5 Perturbations
In the discussions above, we have argued that the superfluid is a gapless SPT is protected by $U(1)$ and $W$. As long as the Hamiltonian (and the boundary conditions) preserve $U(1)$ and the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$, the edge degeneracy is protected. For instance, as verified by DMRG, adding a term $\sum\sigma^{z}_{i-1/2}\sigma^{z}_{i+1/2}$ does not lift the degeneracy. Adding a small perturbation $h_{x}\sum\sigma^{x}_{i+1/2}$ however breaks this symmetry. The edges open up a small gap that scales as $1/L$ and are no longer degenerate (see Appendix E for more details). The situation is different in higher dimensions when the protecting symmetries include higher-form symmetries. Breaking higher-form symmetries explicitly may not lift the edge degeneracy. Furthermore, we could introduce another type of perturbation $h_{z}\sum\sigma^{z}_{i+1/2}$, which preserves the $W$ symmetry, but violates the effective gauge constraint from the large $K$-term. When $h_{z}\ll K$, the gauge constraint will still be effectively enforced at low energy, and exact edge degeneracy remains. Only when $h_{z}<K$ becomes sufficient large, does the degeneracy become lifted, and the edge modes become exponentially localized instead of exactly localized (See Appendix E for more details).
@@
Having considered the 1-D case, we can generalize the analysis to higher dimensions. In this section, we consider the emergent $\mathbb{Z}_{2}$-gauged Bose-Hubbard model on a 2-D square lattice [Fig. 4(a)]:
@@ IV.1 Emergent mixed anomaly
Similar to the 1-D case, the ground state of the system cannot be trivially gapped (i.e., nondegenerate, gapped, and symmetric under both symmetries). This consequence strongly constrains the phase diagram. As we will argue in this work, the critical point inherited from the ordinary superfluid-insulator transition now becomes a generalized DQCP between a gapless SPT phase where $\tilde{U}(1)$ is spontaneously broken and a SET phase where the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry is spontaneously broken [Fig. 4(b)].
@@ IV.2 Gapless SPT phase
To show that the state is SPT, we place it on lattice with open boundaries that preserve the dual ^$\hat{\mathbb{Z}}_{2}$ symmetry $W$. We consider a half-plane geometry with an infinitely long boundary [Fig. 5(a)], where the “rough” boundary has dangling bonds sticking out so that properly modified local gauge constraints Eq. (18) are still well-defined [56]. Then similar to the 1-D case, we can argue that if $W$ is preserved, there will necessarily be a boundary degeneracy as follows. We choose the symmetry generator of $W=\prod_{e\in\gamma}\sigma_{e}^{z}$ to be a Wilson line with one end terminating on the boundary and the other end either extending to infinity or terminating on a boson creation/annihilation operator. With $P$ and $W$ preserved in the bulk, the anti-commutativity of $P$ and $W$ implies a ground state degeneracy, which in this case necessarily comes from the boundary. This defines an SPT phase. In fact, the Higgs condensate $\langle b_{i}^{\dagger}\sigma_{i+1/2}^{z}...\sigma_{j-1/2}^{z}b_{j}\rangle$ can be viewed as an SPT string order parameter here.
@@ IV.3 Excitations in the superfluid phase
We have argued that the superfluid phase is a gapless SPT phase. Here we discuss the excitations in the phase: vortices, Goldstone bosons, and domain walls. Since the $\mathbb{Z}_{2}$-gauging process amounts to projecting out the $\mathbb{Z}_{2}$-charged sector and adding the twisted sector to the theory and the process relates some order parameters to disorder parameters [33, 55], many physical properties of this new gapless SPT phase can be inferred directly from its superfluid parent state. As we have already seen above, the disorder parameter $|\langle\hat{X}_{R}(\theta)\rangle|$ for $\tilde{U}(1)$ is a (fractional) order parameter for the dual ^$\hat{\mathbb{Z}}_{2}$. Similarly, $|\langle b_{i}^{\dagger}(\prod_{e\in\gamma_{ij}}\sigma_{e}^{z})b_{j}\rangle|$ where $\gamma_{ij}$ connects sites $i$ and $j$ [Fig. 5(c)], viewed as a disorder parameter for the dual ^$\hat{\mathbb{Z}}_{2}$, serves as a (fractional) order parameter for $\tilde{U}(1)$.
To guarantee the exactness of the 1-form symmetry, we imposed the zero-flux condition $\prod_{e\in\partial p}\sigma_{e}^{z}=1$, which suppresses all (dynamical) $\pi$-vortices excitations. Equivalently, an open ’t Hooft line which would end on a pair of $\pi$-vortex [Fig. 5(c)] are also suppressed. In the Higgs phase, the condensate phase $\varphi$ is locked to the vortices. The dual 1-form symmetry measures the $\mathbb{Z}_{2}$ winding of condensate phase $\varphi$, which takes values in $n\pi$ for integer $n$ modulo 2, along noncontractible cycles, which is equivalent to insertions of $\pi$-fluxes across noncontractible cycles [see Fig. 5(d)]. We should compare the scenario with that in the “parent” superfluid phase before gauging where there is an emergent 1-form $U(1)$ symmetry in the low energy sector [22, 60, 61]. Charged objects of the emergent $U(1)$ are the winding of $\varphi$, taking values in $2n\pi$ for integer $n$. $2\pi$-vortices explicitly breaks this emergent dual $U(1)$ 1-form symmetry. However, since they are neutral under the dual ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry $W$ after the effective gauging, they do not destroy the exactness of the dual ^$\hat{\mathbb{Z}}_{2}$ 1-form symmetry.